OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 3, -13, and 5 + 4i

OpenStudy (experimentx):

P(x) = (x-3)(x+13)(x-5-4i)(x-5+4i)

OpenStudy (anonymous):

f(x) = x^4 - 8x^3 - 12x^2 + 400x - 1599 f(x) = x^4 - 200x^2 + 800x - 1599 f(x) = x^4 - 98x^2 + 800x - 1599 f(x) = x^4 - 8x^3 + 12x^2 - 400x + 1599

OpenStudy (anonymous):

@experimentX

OpenStudy (debbieg):

Do you understand that for each zero C there must be a factor, (x-C)? And also - any complex root must also have its complex conjugate as a root. So since 5+4i is a root, 5-4i is a root. experimentX used these concepts in setting up the polynomial above, but I just want to make sure you understand them.

OpenStudy (debbieg):

So the set-up is: \[\Large P(x) = (x-3)(x+13)[x-(5-4i)][x-(5+4i)]\]

OpenStudy (debbieg):

Start by dealing with the product involving the complex conjugates. use some neat little algebra tricks to simplify: \(\large [x-(5-4i)][x-(5+4i)]=[x-5+4i][x-5-4i]\) by distribution \(\large =[(x-5)+4i][(x-5)-4i]\) by associativity Now look at that product... it should look familiar, as a form (a+b)(a-b), where a=(x-5) b=4i Take that product and you'll have a 2nd degree polynomial. So you'll be at this point: \[\Large P(x) = (x-3)(x+13)(........)\] Where that last .... is the 2nd degree polynomial obtained above. Then foil the first 2 factors, you'll have 2 polynomials, and then just use distribution.

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Complex_conjugate_root_theorem This states that complex roots always come in pairs a+bi its pair would be a-bi in your case 3, -13, and 5 + 4i AND... 5-4i <- the missing pair :) ---- Hold on i will explain more

OpenStudy (anonymous):

3, -13, 5 + 4i ,5-4i a, b , c , d THe roots can be written like this (x-a) * (x-b) * (x-c) *(x-d) we just plug in (x-2) * (x-(-13)) * (x-c) *(x-d) The first 2 are straitforward. (x-2) * (x-(-13)) * (x-(5 + 4i)) *(x-(5-4i)) c and you just plug in as well. now you expand

OpenStudy (anonymous):

But how do you do that with the i

OpenStudy (anonymous):

lets do that first then (x-(5 + 4i)) *(x-(5-4i)) (x-5-4i)*(x-5+4i) got rid of parenthesis by distributing the - with me so far?

OpenStudy (anonymous):